If a vector A has the following coponents, use trigonometry to find its magnitude and the and the counterclockwise angle it makes with the + x-axis.

Ax= 7.00lb, Ay= 5.00lb

Ax=7.00lb, Ay Please give your answer in the interval [0,360]?

Ax = -21.0m/s , Ay = -34.0m/s

Ax = -21.0m/s , Ay = -34.0m/s

Please give your answer in the interval [0,360]?

Ax = -1600km , Ay = 2500km

Ax = -1600km , Ay = 2500km

Please give your answer in the interval [0,360]?

Ax = 72.0N , Ay = -52.1N

Ax = 72.0N , Ay = -52.1N

Please give your answer in the interval [0,360]?.
Answer #1

TanA=5/7

A=35

TanB=34/21

B=58

TanC=2500/(-1600)

C=122.62

TanD=-52.1/72

D=144.11
Answer #3

The magnitude pf a vector is given by, The angle is given by, e- tan-1 Substitute given values in the equation A-A A 1지 1(71bTherefore, the magnitude of vector A is 39.9 m/s The angle is given by, 0tan--34 m/s 21 m/s - 58.29 The counter clockwise angSubstitute given values in the equation A-JA;+A; A72N) (-52.1N) -88.87 N Therefore, the magnitude of vector A is 88.87 N The